Shift summary day 25-Sept.-1999

Nickolas Solomey (solomey@hep.uchicago.edu)
Sat, 25 Sep 1999 15:20:17 -0500

Shift Summary for day shift, 25-Sept.-1999

On Shift: Rick Ford, and Nick

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0h 0min

Accounting of lost time:

NM2EU power supply failure from Owl shift. Please also note that to replay
this supply the crew had to disconnect and move then reinstall the power
supplies for NM2S2 and NM4AN-A These should be watched carefully when we
come back up since it would be nice to have the same magnet setting for
both the analysis magnet and the Cascade-0 polarization.

========================================================================

Runs ended this shift: None

Tapes brought to FCC this shift: None

Accesses during shift: None.

Overview of shift: Waiting for the power supply to be swapped.

========================================================================

Description of Shift: Bob T. Debugged trigger 16, no success.
Jasson Hamm worked on HCC stuff, that is all he would mumble when asked.

At 1:30 we just received another 3 hr estimate from the power supply
techs. It seems that the old NM2EU supply had a modified transducter which
allowed more stable operation, which must be installed in the replacement.
this added an extra hour.

========================================================================
Follow up on next shift:

Bob T. passed the word that we should make sure when beam comes
back up that we do a muon run first.